Could someone tell me how to do this?
A trapezoid has height 8 and area 28. What are the lenghts of it's bases?

Answer by checkley71(8403)

THE AREA OF A TRAPEZOID IS (BASE1+BASE2)/2*HEIGHT THEREFORE (BASE1+BASE2)/2*8=28 OR (BASE1+BASE2)*4=28 OR BASE1+BASE2=28/4 OR BASE1+BASE2=7